Problem 3-2A (Part Level Submission) The Concord Corporation opened for business on May 1, 2019. Its trial balance before adjustment on May 31 is as follows. Concord Corporation Trial Balance May 31, 2019 Debit Account Number 101 126 130 140 141 149 201 208 275 311 429 610 726 732 Credit Cash Supplies Prepaid Insurance Land Buildings Equipment Accounts Payable Unearned Rent Revenue Mortgage Payable Common Stock Rent Revenue Advertising Expense Salaries and Wages Expense Utilities Expense $3,500 2,150 1,800 12,000 58,800 15,600 11,900 3,400 40,000 35,700 7,550 650 3,200 850 $98,550 In addition to those accounts listed on the trial balance, the chart of accounts for Concord Corporation also contains the following accounts and account numbers: No. 142 Accumulated Depreciation-Buildings, No. 150 Accumulated Depreciation-Equipment, No. 212 Salaries and Wages Payable, No. 230 Interest Payable, No. 619 Depreclation Expense, No. 631 Supplies Expense, No. 718 Interest Expense, and Noin addition to those accounts listed on the trial balance, the chart of accounts for Concord Corporation also contains the following accounts and account numbers: No. 142 Accumulated Depreciation-Buildings, No. 150 Accumulated Depreciation-Equipment, No. 212 Salaries and Wages Payable, No. 230 Interest Payable, No. 619 Depreciation Expense, No. 631 Supplies Expense, No. 718 Interest Expense, and No. 722 Insurance Expense. Other data 1. Prepaid insurance is a 1-year policy starting May 1, 2019 2. A count of supplies shows $800 of unused supplies on May 31. 3 Annual depreciation is $2,940 on the buildings and $1,560 on equipment. 4 The mortgage interest rate is 12%. (The mortgage was taken out on May 1.) 5. Two-thirds of the unearned rent revenue has been earned. 6. Salaries of $700 are accrued and unpaid at May 31. Journalize the adjusting entries on May 31. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. Round answers to 0 decimal places, e.g. 5,275.) No. Date Account Titles and Explanation 1. May 31 Insurance Expense Debit Credit Prepaid Insurance 2. May 31 Supplies Expense Supplies 3. May 31 Depreciation Expense Depreclation-Buildings Accumulated Deoreciation-Eouioment4. May 31 Interest Expense Interest Payable S. May 31 Unearned Rent Revenue Rent Revenue 6. May 31 Salaries and Wages Expense Salarles and Wages Payable Attempts: 0 of 3 used SAVE FOR LATER SUBMIT ANSWER
Answer #1
1 May 31 Insurance expense 150 =1800/12
        Prepaid insurance 150
2 May 31 Supplies expense 1350 =2150-800
       Supplies 1350
3 May 31 Depreciation expense 375
       Accumulated Depreciation-Buildings 245 =2940/12
       Accumulated Depreciation-Equipment 130 =1560/12
4 May 31 Interest expense 400 =40000*12%/12
       Interest payable 400
5 May 31 Unearned rent revenue 2267 =3400*2/3
       Rent revenue 2267
6 May 31 Salaries and wages expense 700
      Salaries and wages payable 700